Im in business math,I haven't been in school for a while,I'm so sad I'm not doing well,.
#3.FIND THE EQUIVALANT EARNINGS FOR EACH PART:
a. Annual earnings= $125,678.60
Weely earning=
Biweekly earnings=
Monthly earnings=
Semimonthly earnings=
I also don't understand part,base, & rate. Also page 4 on my homework,I'm nervous, please help me 😞
Thank You, Dana Howie

Answer by mananth(16949)

a. Annual earnings= $125,678.60
Weely earning=
weeks in year = 52
earnings /week = 125,678.60/52= $2,416.90

Biweekly earnings=2416.90/2= $1208.45
Monthly earnings=
12 months
$125,678.60/12 =$10,473.22
Semimonthly earnings=
earnings per week = $2416.90
semimonthly = 2416.90 * 2= $4,833.80